# Cyclotomic Factors and LRS-Degeneracy

### v3

@@ 4.1.3 Algorithm
The inner “refinement” algorithm takes 3 inputs: the polynomial $f$, the evaluation point $\beta\in{\mathbb{Q}}_{>1}$, and a list of candidate indexes. It produces a refined list of candidate indexes, which might be empty. If $f$ is known to be palindromic, we simply set $N\leftarrow f(\beta)_{num}$ in Step 3.
@@
13:  Set $N\leftarrow N/\gcd(N,g^{\infty})$
@@
The outer algorithm takes just the polynomial $f$ as input; it could also perform the preprocessing steps mentioned earlier. It checks explicitly for the factors $\Phi_{1}$ and $\Phi_{2}$, so that calls to the inner algorithm are not complicated by the exceptions in Zsigmondy’s theorem. It always chooses $2$ as the first evaluation point (for reasons of computational speed); thereafter it picks random small rationals. The randomness serves to preclude the possibility of constructing a small input which causes the algorithm to perform poorly — this is standard practice in randomized algorithms.
@@
4:Set $L_{12}\leftarrow[\,]$. If f(1)=0f(1){=}0, append 11 to $L_{12}$. If f(1)=0f(-1){=}0, append 22 to $L_{12}$.
5:Set $L$ to be a list containing all $k\in{\mathbb{N}}_{>2}$ for which $\phi(k)\leq\deg(f)$
6:Set $L\leftarrow\textbf{RefineCandidates}(f,2,L)$ $\triangleright$ see Algorithm 2
@@
9:  Set $L_{new}\leftarrow\textbf{RefineCandidates}(f,\beta,L)$ $\triangleright$ see Algorithm 2
@@
11:  Set $L\leftarrow L_{new}$, and perform another iteration
@@
If the input polynomial is monic and of even degree, we can perform, before Step 5, an initial cyclotomicity test (from Section 3) as this is very quick. The initial refinement at $\beta=2$ was quite effective in our tests: i.e. only few false positives survived. As written, Algorithm 3 has difficulty excluding the indexes $3,4$ and $6$; so in Step 10, the first time the condition $L_{new}=L$ is satisfied, we perform one more iteration but using a $\beta$ value chosen so that each of $\Phi_{k}(\beta)_{num}$ for $k=3,4,6$ has a prime factor larger than 10 00010\,000, e.g. suitable $\beta$ values are $\frac{117}{98}$, $\frac{133}{18}$ or $\frac{169}{6}$.
@@
Let $f=\sum_{j=0}^{d}a_{j}x^{j}$ be non-constant. Let $p$ be a prime which divides $a_{0},a_{1},\ldots,a_{r}$ and $a_{d},a_{d-1},\ldots,a_{d-s}$; note we do not need to compute $p$, it suffices to verify that $\gcd(a_{0},a_{1},\ldots,a_{r},\;a_{d},a_{d-1},\ldots,a_{d-s})\neq 1$. Suppose $\Phi_{k}\mid f$, then $\psi_{p}(\Phi_{k})\mid x^{-(r{+}1)}\,\psi_{p}(f)$ where $\psi_{p}$ is reduction modulo $p$. Then we obtain the better bound for $\phi(k)=\deg(\Phi_{k})\leq\deg(f)-(r{+}s{+}2)$ in Step 5 of Algorithm 3. If $f$ is palindromic then we have $r=s$.
@@
We use a simple technique for generating “random rationals”. We regard the rationals in ${\mathbb{Q}}_{>1}$ as being ordered lexicographically via the mapping $p/q\mapsto(p,q)$ with $\gcd(p,q)=1$. A new rational was chosen by jumping forward along this progression by a random amount; by always jumping forward, we avoid generating the same value twice.
@@ 5.1 Some properties of LRS-degenerate polynomials
  • Let $f\in{\mathbb{Z}}[x]$ be divisible by $x^{d}$ for some $d\in{\mathbb{N}}$ then $f$ is $k$-LRS-degenerate iff $f/x^{d}$ is.
@@ 5.2 Preprocessing prior to testing LRS-degeneracy
  • Optionally, we may also simplify the coefficients using Algorithm 4 from Section 5.2.1.
@@ 5.2.1 Preprocessing: coefficient reduction
4:Compute $g\leftarrow\gcd(a_{1},a_{2},\ldots,a_{d})$ $\triangleright$ $a_{0}$ is deliberately excluded
@@
8:  Compute $m\leftarrow\min\{\lfloor\mu_{p}(a_{j})/j\rfloor\mid j=1,2,\ldots,d\}$
@@
10:  Set $\lambda\leftarrow\lambda/p^{m}$
@@ 5.4 Recalling CDM Algorithm 1
The first algorithm in [3] computes $R_{f}=\mathop{\rm res}_{y}(f(y),f(xy))/(x-1)^{d}$ where $d=\deg(f)$. So $R_{f}$ is a palindromic polynomial of degree $d^{2}-d$, since we assume that $f$ is square-free. It then determines whether $R_{f}$ has any cyclotomic factors (with index $\geq 2$), since $f$ is $k$-LRS-degenerate iff $\Phi_{k}(x)$ divides $R_{f}$. In practice, they observe that this method becomes rather slow for $\deg(f)>25$ because the resultant computation is costly. Also they sought cyclotomic factors by first computing a factorization into irreducibles, which can be slow in some cases. Our Algorithm 3 or that from [2] would make the detection of cyclotomic factors much faster, but would not reduce the cost of computing the resultant. Our Algorithm 3 produces also the orders of LRS-degeneracy.
@@ 5.6 The modular algorithm
The loop at Step 6 in Algorithm 5 performs 3 iterations because that gave a good compromise between speed of computation and exclusion of false positives.
@@ 5.8 Test for $k$-LRS-degeneracy
3:Set $K$ to be the list of candidate orders to test $\triangleright$ see below
4:Set $L\leftarrow[\,]$, it will contain all “probable” orders detected
@@ 5.9 Which $k$ values to try?
In Step 3 of Algorithm 5 we need an initial set of candidate orders to test. Recall that we assume that $f$ is square-free in ${\mathbb{Z}}[x]$; we write $d=\deg(f)$. One possible set comprises all $k$ such that $\phi(k)\leq d^{2}-d$, because if $f$ is $k$-LRS-degenerate then $\Phi_{k}(x)$ divides $\mathop{\rm res}_{y}(f(xy),f(y))/(x-1)^{d}$ which has degree $d^{2}-d$. Below we explain how to reduce this set somewhat.
@@
3:Set $D\leftarrow[\,]$, this will be our list of values which $\phi(k)$ must divide
@@ 5.9.1 Conjectured short-cut
Let $f=\Phi_{3}\Phi_{5}$, so $\deg(f)=6$. Then $f$ is obviously $k$-LRS-degenerate for $k=3$ and $k=5$. In fact, $f$ is also $15$-LRS-degenerate, yet $\deg(\Phi_{15})=8$. So if we want to find all orders $k$ then the conjectured limit does not apply.

### v4

@@ 4.1.3 Algorithm
The inner “refinement” algorithm takes 3 inputs: the polynomial $f$, the evaluation point $\beta\in{\mathbb{Q}}_{>1}$, and a list of candidate indexes. It produces a refined list of candidate indexes, which might be empty.
@@
13:  $N\leftarrow N/\gcd(N,g^{\infty})$ $\triangleright$ See Remark 4.2
@@
Remark 4.2.
In Algorithm 3 (RefineCandidates) at Step 10 we write $N/\gcd(N,g^{\infty})$ to mean $\lim_{s\to\infty}N/\gcd(N,g^{s})$; it is the largest factor of $N$ coprime to $g$.
The outer algorithm takes just the polynomial $f$ as input; it could also perform the preprocessing steps mentioned earlier. It checks explicitly for the factors $\Phi_{1}$ and $\Phi_{2}$, so that calls to the inner algorithm are not complicated by the exceptions in Zsigmondy’s theorem. It always chooses $2$ as the first evaluation point (for reasons of computational speed); thereafter it picks random small rationals. The randomness serves to preclude the possibility of constructing a small input which consistently causes the algorithm to perform poorly — this is standard practice in randomized algorithms.
@@
[…]
@@
9:  $L_{new}\leftarrow\textbf{RefineCandidates}(f,\beta,L)$ $\triangleright$ see Algorithm 3
@@
11:  $L\leftarrow L_{new}$, and perform another iteration
@@
If the input polynomial is monic and of even degree, we can perform, before Step 5, an initial cyclotomicity test (from Section 3) as this is very quick. The initial refinement at $\beta=2$ was quite effective in our tests: i.e. only few false positives survived. As written, Algorithm 4 (FindCycloFactors) has difficulty excluding the indexes $3,4$ and $6$; so in Step 10, the first time the condition $L_{new}=L$ is satisfied, we perform one more iteration but using a $\beta$ value chosen so that each of $\Phi_{k}(\beta)_{num}$ for $k=3,4,6$ has a prime factor larger than 10 00010\,000, e.g. suitable $\beta$ values are $\frac{117}{98}$, $\frac{133}{18}$ or $\frac{169}{6}$.
@@
Let $f=\sum_{j=0}^{d}a_{j}x^{j}$ be non-constant. Let $p$ be a prime which divides $a_{0},a_{1},\ldots,a_{r}$ and $a_{d},a_{d-1},\ldots,a_{d-s}$; note we do not need to compute $p$, it suffices to verify that $\gcd(a_{0},a_{1},\ldots,a_{r},\;a_{d},a_{d-1},\ldots,a_{d-s})\neq 1$. Suppose $\Phi_{k}\mid f$, then $\psi_{p}(\Phi_{k})\mid x^{-(r{+}1)}\,\psi_{p}(f)$ where $\psi_{p}$ is reduction modulo $p$. Then we obtain the better bound for $\phi(k)=\deg(\Phi_{k})\leq\deg(f)-(r{+}s{+}2)$ in Step 5 of Algorithm 4 (FindCycloFactors). If $f$ is palindromic then we have $r=s$.
@@
We use a simple technique for generating “random rationals”. We regard the rationals in ${\mathbb{Q}}_{>1}$ as being ordered lexicographically via the mapping $p/q\mapsto(p,q)$ with $\gcd(p,q)=1$. A new rational is chosen by jumping forward along this progression by a random amount; by always jumping forward, we avoid generating the same value twice.
@@ 5.1 Some properties of LRS-degenerate polynomials
  • Let $f\in{\mathbb{Z}}[x]$ be divisible by $x^{d}$ for some exponent $d\in{\mathbb{N}}$ then $f$ is $k$-LRS-degenerate iff $f/x^{d}$ is.
@@ 5.2 Preprocessing prior to testing LRS-degeneracy
  • Optionally, we may also simplify the coefficients using Algorithm 5 (ReduceCoefficients) from Section 5.2.1.
@@ 5.2.1 Preprocessing: coefficient reduction
[…]
@@
8:  $m\leftarrow\min\{\lfloor\mu_{p}(a_{j})/j\rfloor\mid j=1,2,\ldots,d\}$
@@
10:  $\lambda\leftarrow\lambda/p^{m}$
@@ 5.4 Recalling CDM Algorithm 1
The first algorithm in [9] computes $R_{f}=\mathop{\rm res}_{y}(f(y),f(xy))/(x-1)^{d}$ where $d=\deg(f)$. So $R_{f}$ is a palindromic polynomial of degree $d^{2}-d$, since we assume that $f$ is square-free. It then determines whether $R_{f}$ has any cyclotomic factors (with index $\geq 2$), since $f$ is $k$-LRS-degenerate iff $\Phi_{k}(x)$ divides $R_{f}$. In practice, they observe that this method becomes rather slow for $\deg(f)>25$ because the resultant computation is costly. Also they sought cyclotomic factors by first computing a factorization into irreducibles, which can be slow in some cases. Our Algorithm 4 (FindCycloFactors) or that from [14] would make the detection of cyclotomic factors much faster, but would not reduce the cost of computing the resultant. Our Algorithm 4 (FindCycloFactors) produces also the orders of LRS-degeneracy.
@@ 5.6 The modular algorithm
In Algorithm 6 (LRSDegeneracyOrders) the loop from Step 6 to Step 13 performs 33 iterations because that gave a good compromise between speed of computation and exclusion of false positives.
@@ 5.8 Test for $k$-LRS-degeneracy
[…]
@@ 5.9 Which $k$ values to try?
In Step 3 of Algorithm 6 (LRSDegeneracyOrders) we need an initial set of candidate orders to test. Recall that we assume that $f$ is square-free in ${\mathbb{Z}}[x]$; we write $d=\deg(f)$. One possible set comprises all $k$ such that $\phi(k)\leq d^{2}-d$, because if $f$ is $k$-LRS-degenerate then $\Phi_{k}(x)$ divides $\mathop{\rm res}_{y}(f(xy),f(y))/(x-1)^{d}$ which has degree $d^{2}-d$. Below we explain how to reduce this set somewhat.
@@
[…]
@@ 5.9.1 Conjectured short-cut
Let $f=\Phi_{3}\Phi_{5}$, so $\deg(f)=6$. Then $f$ is obviously $k$-LRS-degenerate for $k=3$ and $k=5$. In fact, $f$ is also $15$-LRS-degenerate, yet $\phi(15)=8>\deg(f)$. So if we want to find all orders $k$ then the conjectured limit does not apply.